# Multi-fingered Robotic Hand Grasping in Cluttered Environments through Hand-Object Contact Semantic Mapping

### v2

@@ Abstract
The deep learning models has significantly advanced dexterous manipulation techniques for multi-fingered hand grasping. However, the contact information-guided grasping in cluttered environments remains largely underexplored. To address this gap, we have developed a method for generating multi-fingered hand grasp samples in cluttered settings through contact semantic map. We introduce a contact semantic conditional variational autoencoder network (CoSe-CVAE) for creating comprehensive contact semantic map from object point cloud. We utilize grasp detection method to estimate hand grasp poses from the contact semantic map. Finally, an unified grasp evaluation model is designed to assess grasp quality and collision probability, substantially improving the reliability of identifying optimal grasps in cluttered scenarios. Our grasp generation method has demonstrated remarkable success, outperforming state-of-the-art methods by at least 4.65% with 81.0% average grasping success rate in real-world single-object environment and 75.3% grasping success rate in cluttered scenes. We also proposed the multi-modal multi-fingered grasping dataset generation method. Our multi-fingered hand grasping dataset outperforms previous datasets in scene diversity, modality diversity. The dataset, code and supplementary materials can be found at https://sites.google.com/view/ffh-cluttered-grasping.
@@ I Introduction
Utilizing a multi-fingered hand for robotic grasping enhances both ease and flexibility, primarily through the increased number of contact points available for interaction. Recent advancements in multi-fingered robotic grasping research [1, 2, 3, 4, 5, 6, 7] have focused on leveraging hand-object contact information to guide the generation of grasping strategies. By utilizing contact points and contact distance maps, respectively, UniGrasp [1] and GenDexGrasp [2] guide the grasp generation process, thereby enhancing the generalization of grasping strategies across different robotic hands. However, the contact points generated by UniGrasp [1] are sparse, limiting the diversity of the generated grasps. The contact distance maps lack semantic information, which leads to instability in the grasping samples generated under their guidance. To solve these limitations of contact information-guided grasp generation methods, we propose the CoSe-CVAE to generate contact semantic maps for assisting grasp generation.
In addition to contact information-guided grasp generation methods, recent advancements in multi-fingered hand grasping have enabled the grasp generation from visual data through deep learning methods [8, 9, 10], generation model [11, 12]. However, real-world scenarios are often cluttered, making grasp planning for multi-fingered hands particularly challenging. Additionally, the high degrees of freedom in multi-fingered hands further complicate the generation and evaluation of collision-free, high-quality grasps. Despite numerous studies [13, 8, 14, 15, 16, 17, 18, 19, 20] exploring the grasping in cluttered scenes using various end-effectors, their ability to reliably evaluate multi-fingered hand grasping in such environments remains limited. HGC-Net [8] remains unstable when evaluating grasps colliding with the surroundings. Moreover, these methods are often tailored to specific robotic hand designs, lacking the flexibility to generalize across different multi-fingered hands. To address mentioned limitations, we introduce a grasp evaluator to estimate grasp quality and collision probability, leveraging point cloud data from both the object and the robotic hand.
@@
Figure 1: Employing CoSe-CVAE, the contact semantic maps are derived from object point clouds. Grasp detection leverages contact prior information for estimating grasp poses. Subsequent grasp evaluation assesses both grasping qualities with collision awareness to identify the optimal grasp in cluttered settings. (Blue: grasp candidates colliding with the surroundings, Red: negative grasp candidates, Green: Positive grasp candidates. )
Our research introduces a multi-fingered hand grasping generation pipeline using contact semantic map in cluttered scenes. This pipeline includes the CoSe-CVAE, a grasp detection method and the grasp evaluation model. CoSe-CVAE is designed to generate contact semantic maps, where the semantic information indicates which fingers are in contact with the object. Furthermore, the grasp poses are estimated based on predicted contact semantic maps and optimal grasp is selected based on grasp qualities from grasp evaluation model. Our main contributions are as follows:
@@
    We propose a contact semantic conditional variational autoencoder network (CoSe-CVAE) that generates multi-fingered grasping contact semantic maps from object point clouds. CoSe-CVAE generates richer, more diverse contact point maps with semantic information, enabling more stable and reliable grasp generation guided by contact information.
@@
    We introduce a grasp evaluation network estimating grasp scores by analysing the partial scene point cloud and hand geometric features based on PointNet++ [29]. The network is capable of evaluating grasping in cluttered scenes for both known and unknown multi-fingered hands. Our method outperforms SOTA approaches [8, 2, 1] in average grasp success rate by at least 4.65%.
@@
    We integrate a pipeline for generating a multi-modal multi-fingered grasping dataset in cluttered environments, based on DexGraspNet [28]. Its scene complexity and number of modalities surpass those of previous five-finger hand datasets.
@@ II-A Multi-fingered Robotic Hand Grasping in Cluttered Environments
Grasping in cluttered environments using multi-fingered robotic hands presents a significant challenge due to their high degrees of freedom and the complex collision dynamics with surrounding objects. Compared to two-jaw grippers, multi-fingered robotic hands can establish more contact points, allowing them to grasp more complex objects with enhanced dexterity. Although there has been extensive research [30, 18, 31] on grasping in cluttered environments with two-jaw gripper, studies on multi-fingered robotic hand grasping in such environments remain limited [8, 14, 15, 32]. Currently, datasets for multi-fingered robotic hand grasping in cluttered environments are severely limited. We provide an overview of existing multi-fingered hand grasping datasets and their available modalities, as summarized in Tab. I. However, there are no dataset that includes cluttered scenes while capturing all relevant multi-modal information. To date, no studies have utilized contact information to guide grasp generation in cluttered environments with multi-fingered robotic hands, and no corresponding datasets have been developed. To address this gap, we extended the existing grasp generation pipeline [28] to cluttered environments, producing contact semantic maps. Building on this, we propose a novel contact information-guided method for multi-fingered robotic hand grasp generation in cluttered scenes.
@@ II-B Contact Information-guided Grasping Generation
Hand-object representations are widely used in various domains: they are crucial for generating plausible hand poses [4, 5, 6], formulating generalized representations for diverse end-effectors  [1, 2], and bridging the gap between human and robotic hand representations  [33, 7]. Various types of contact representations are employed, including contact touch code [33], contact distance or points [34, 3, 1], contact semantic map [22, 21]. UniGrasp [1] introduced a generalized model that sequentially generates contact points. Generative models, renowned for their diversity and generative capabilities, have been increasingly applied in the field of grasp generation. Notable examples include the application of CVAE [2, 35, 36] and diffusion model [37]. GenDexGrasp [2] employed a generative model to generate contact distance maps from object point clouds. However, these methods do not address how to generate contact semantic map from scene point cloud and generate stable multi-fingered hand grasps using contact semantic map in complex environments. Additionally, we found that grasps generated using contact distance maps lacked stability due to the absence of semantic information. To address these limitations, we propose a novel generative model, named CoSe-CVAE, which generates contact semantic maps from object point clouds and incorporates the grasp generation pipeline for cluttered environments.
@@ II-C Grasping Evaluation in Cluttered Scenes
Concerning collision-free grasp detection in cluttered environments, extensive research has been conducted on employing neural networks to predict collision-free grasp samples from visual data, particularly in the context of two-fingered grasping setups [30, 16, 8, 17, 18, 19, 38, 20]. Our previous work [30] developed a cable grasping neural network (CG-CNN) aiming at estimating grasp quality for cables in cluttered scenes, using implicit collision representations. However, multi-fingered hands, with their additional joints, pose greater challenges in learning implicit collision representations. Previous evaluators [39] were often designed for a specific robotic hand and lacked the ability to generalize across different hand types. To address this complexity and identify optimal grasp candidates in cluttered environments, we develop an unified grasp evaluation model that estimate grasp scores with collision awareness from partial scene point cloud and the point cloud of robotic hand, while being adaptable to various robotic hands.
@@ III-A Problem Statement
We utilize contact semantic map $\Omega\in\mathbb{R}^{2048\times(n+1)}$ with $2048$ points to represent the contact points between $n$ fingers of robotic hand and grasped object and points without contacts, as shown in Fig. 2 (a). Generative model CoSe-CVAE $f$ is able to estimate $N$ contact semantic maps from object point cloud $O$, as introduced in Sec. III-C. Grasp detection $F$ is presented in Sec. III-D to estimate grasp candidates based on contact semantic maps. For estimating the optimal grasp candidate $g_{\rm optimal}$ from a cluttered scene’s point cloud, grasp evaluation network $\Psi$ infers grasp qualities $q$ from partial scene point cloud $P$ and sampled hand point cloud $H$, as described in Sec. III-E. The partial scene point cloud is obtained through filtering the original scene points using a cylindrical region in the robotic hand’s frame, defined by a radius $r$ and height $h$, as depicted in Fig. 2 (c). The optimal grasp pose $g_{\rm optimal}$ is selected based on these inferences. The pipeline is summarized in Eq. 1.
@@ III-B Multi-Modal Multi-fingered Hand Grasping Dataset Generation
1:Input: Object database, Number of sampling grasp poses $M$, Number of objects in scene $m$
2:Output: Set of multi-fingered robotic hand grasping candidates with grasp pose $g$, grasp quality $Q_{1}$, collision score $Q_{2}$, contact semantic map $\Omega$
3:Procedure 1: Estimate grasp pose $g$, grasp qualities $Q_{1}$ based on [28], contact distance map based on [2].
4:Procedure 2: Contact semantic mapping estimation as detailed in III-B1
5:Procedure 3: Cluttered scene generation and collision evaluation
6:repeat
7:     Build cluttered scene with $m$ objects based on collision checking among objects [18].
8:     Estimate collision state between $j$ grasping candidates and corresponding surrounding objects to obtain collision scores $\bigcup\limits_{j=0}^{M-1}{Q}_{2j}$. The collision score of $j$th candidate is denoted by $Q_{2j}$.
9:until Finish collision checking for all multi-fingered robotic grasping samples for cluttered scenes.
@@ III-B1 Contact Semantic Map Estimation
[…]
@@
Figure 4: Contact Semantic Map Generation and Grasping Detection.
@@ III-D Grasping Detection from Contact Semantic Maps
Using the generated contact semantic maps and the surface point clouds of each finger of the robotic hand, we utilize the correspondence point matching algorithm [41] to estimate the initial wrist pose. Inspired by GenDexGrasp [2] and UniGrasp [1], we optimize the wrist poses and finger tip positions by minimizing energy loss and the joint angles of the manipulator are calculated by differential inverse kinematics library mink [42]. The energy loss function $e$ is formulated as follows:
@@ III-E Grasp Evaluation Model
To identify optimal grasp in cluttered environments, we introduce an unified grasp evaluation model to quantify grasp quality, denoted as $q$. As illustrated in Fig. 5, our network architecture integrates inputs composed of hand surface points $P$ and a partial scene point cloud $H$ surrounding the target grasped object. The partial point cloud is captured relative to the hand’s local frame. The point clouds are processed through two PointNet++ encoders, which extract latent spatial features of the hand and the scene point clouds. These latent features are concatenated and passed through a PointNet++ decoder to predict grasp classification scores. The classification includes three categories: Class 0 (grasp candidates under collision condition), Class 1 (negative grasp candidates), and Class 2 (positive grasp candidates). We employ a multi-class classification loss to guide the network’s training. Specifically, we utilize the categorical cross-entropy loss, defined as:
@@ IV-B Multi-fingered Robotic Hand Grasping Dataset
Our dataset includes 1,521 household models from our previous work, as well as objects from the AffordPose dataset [46]. We generated over 2,000 scenes containing cluttered scene point clouds, collision scores, grasp quality, contact semantic maps, and grasp type information, contact distance map. We predict grasp types by voting based on the contact points and the object affordance label from [46], where the grasp type is determined by the affordance with the highest number of votes. The affordance labels include handle grasping, enveloping grasp, pouring, pressing, cutting, and twisting. In this work, we did not explore grasp type information in grasp generation. Dataset examples and grasp types are shown in the Fig. 3 and Fig. 7.
@@ IV-C Model Training
Figure 8: Inference results for scene with single object and cluttered scenes. (a) Contact points and grasp candidate based on UniGrasp [1]. (b) Estimated Contact distance map and grasp pose from GenDexGrasp [2]. (c) Contact semantic map and grasp candidate from our methods. d) Grasp candidates based on HGC-Net [8]. (d) Positive grasp candidates based on our methods.
IV-D Qualitative Study with SOTA methods

### v3

# ContactDexNet: Multi-fingered Robotic Hand Grasping in Cluttered Environments through Hand-Object Contact Semantic Mapping

@@ Abstract
The deep learning models has significantly advanced dexterous manipulation techniques for multi-fingered hand grasping. However, the contact information-guided grasping in cluttered environments remains largely underexplored. To address this gap, we have developed ContactDexNet, a method for generating multi-fingered hand grasp samples in cluttered settings through contact semantic map. We introduce a contact semantic conditional variational autoencoder network (CoSe-CVAE) for creating comprehensive contact semantic map from object point cloud. We utilize grasp detection method to estimate hand grasp poses from the contact semantic map. Finally, an unified grasp evaluation model PointNetGPD++ is designed to assess grasp quality and collision probability, substantially improving the reliability of identifying optimal grasps in cluttered scenarios. Our grasp generation method has demonstrated remarkable success, outperforming state-of-the-art (SOTA) methods by at least 4.7%, with 81.0% average grasping success rate in real-world single-object grasping using a known hand, and by at least 9.0% when using an unknown hand. Moreover, in cluttered scenes, our method attains a 76.7% success rate, outperforming the SOTA method by 6.3%. We also proposed the multi-modal multi-fingered grasping dataset generation method. Our multi-fingered hand grasping dataset outperforms previous datasets in scene diversity, modality diversity. More details and supplementary materials can be found at https://sites.google.com/view/contact-dexnet.
@@ I Introduction
Recent advancements in multi-fingered robotic grasping research [1, 2] and human grasp generation [3, 4, 5] have focused on leveraging hand-object contact information to guide the generation of grasping strategies. Specifically, contact information such as contact points from UniGrasp [1] and contact distance map from GenDexGrasp [2] has been shown to enhance the generalizability of grasp generation for previously unknown robotic hands. Additionally, contact maps can facilitate the synthesis of functional human grasp postures [6]. However, existing approaches still face challenges in contact information-guided robotic grasp generation, including low robustness due to sparse contact points [1], semantic ambiguity arising from the absence of contact semantic information [2], high model complexity [3], and infeasible grasp poses resulting from the structural differences between human and robotic hands [3]. To address these limitations, we propose the CoSe-CVAE to generate contact semantic maps to improve multi-fingered robotic hand grasp generation.
@@
Figure 1: Employing CoSe-CVAE, the contact semantic maps are derived from object point clouds. Grasp detection leverages contact prior information for estimating grasp poses. Subsequent grasp evaluation model PointNetGPD++ assesses both grasping qualities with collision awareness to identify the optimal grasp in cluttered settings. (Blue: grasp candidates colliding with the surroundings, Red: negative grasp candidates, Green: Positive grasp candidates. )
Real-world robotic grasping scenarios are often cluttered, making grasp planning for multi-fingered hands particularly challenging. While numerous studies have explored grasping in such scenarios using two-jaw grippers [7, 8, 9] and multi-fingered robotic hands [10, 11, 12, 13, 14], existing multi-fingered robotic grasping methods still struggle to achieve robust and reliable performance. Specifically, many approaches fail to effectively evaluate and execute accurate multi-fingered grasps when potential collisions with surrounding objects exist [11]. Grasp prediction errors may cause premature finger contact, pushing the object away instead of achieving a stable grasp. This issue can be mitigated by incorporating contact information between perception and grasp execution. Furthermore, many existing methods are designed for specific robotic hand models, limiting their adaptability across different multi-fingered hands [11]. This lack of generalizability increases the cost of data collection and model training for grasping with unknown robotic hands. Inspired by our previous work [15], we introduce a generalizable grasp evaluator PointNetGPD++ that estimates grasp quality and collision probability. Designed for broad applicability, our approach enhances adaptability across different robotic hand models and diverse grasping scenarios.
@@
Our research introduces a contact information-guided multi-fingered robotic grasp generation pipeline in cluttered scenes, leveraging contact semantic maps to enhance grasp quality and adaptability. This pipeline includes the CoSe-CVAE, a grasp detection method and a generalizable grasp evaluation model (PointNetGPD++). CoSe-CVAE is designed to generate contact semantic maps, where the semantic information indicates which fingers are in contact with the object. Furthermore, the grasp poses are estimated based on predicted contact semantic maps and optimal grasp is selected based on grasp qualities from grasp evaluation model. Our main contributions are as follows:
@@
    We propose a contact semantic conditional variational autoencoder network (CoSe-CVAE) that generates multi-fingered grasping contact semantic maps from object point clouds. CoSe-CVAE generates richer, more diverse contact point maps with semantic information, enabling more stable and reliable grasp generation guided by contact information. It improves the grasping success rate using known and unknown hands by at least 4.7% and 9.0%.
@@
    We introduce a generalizable grasp evaluation network (PointNetGPD++) estimating grasp scores by analysing the partial scene point cloud and hand geometric features based on PointNet++ [26]. The network is capable of evaluating grasping in cluttered scenes for both known and unknown multi-fingered hands. Our method outperforms SOTA approaches [11, 2, 1, 3] in average grasp success rate by at least 4.7% for grasping from single-object scenes and by 6.3% for grasping from cluttered scenes.
@@
    We integrate a pipeline for generating a multi-modal multi-fingered grasping dataset in cluttered environments, based on DexGraspNet [24]. Compared to previous multi-fingered hand datasets, our dataset includes more complex scenes, a greater number of modalities, and newly introduced contact semantic maps, which enhance grasp representation. Moreover, these maps improve transferability across different robotic hands, enabling broader applicability.
@@ II-A Multi-fingered Robotic Hand Grasping in Cluttered Environments
Grasping in cluttered environments using multi-fingered robotic hands presents a significant challenge due to their high degrees of freedom and the complex collision dynamics with surrounding objects. Although there has been extensive research on grasping in cluttered environments with two-jaw gripper [27, 7] and multi-fingered hand grasping from single-object scenes [28, 29, 30, 31, 32, 33], studies on multi-fingered robotic hand grasping in such environments remain limited [10, 11, 14, 20, 13, 34]. Currently, datasets for multi-fingered robotic hand grasping in cluttered environments are severely limited. We provide an overview of existing multi-fingered hand grasping datasets and their available modalities, as summarized in Tab. I. However, there are no dataset that includes cluttered scenes while capturing all relevant multi-modal information. To date, no studies have utilized contact information to guide grasp generation in cluttered environments with multi-fingered robotic hands, and no corresponding datasets have been developed. To address this gap, we extended the existing grasp generation pipeline [24] to cluttered environments, producing contact semantic maps.
@@ II-B Contact Information-guided Grasping Generation
Hand-object representations are widely used in various domains: they are crucial for generating plausible hand poses [4, 5, 35], formulating generalized representations for diverse end-effectors [1, 2], and bridging the gap between human and robotic hand representations  [36, 37]. Various types of contact representations are employed, including contact touch code [36], contact distance or points [38, 39, 1], contact semantic map [3, 16]. UniGrasp [1] introduced a generalized model that sequentially generates contact points. Generative models, renowned for their diversity and generative capabilities, have been increasingly applied in the field of grasp generation [2, 40, 41, 42]. GenDexGrasp [2] employed a generative model to generate contact distance maps from object point clouds. However, we found that grasps generated using contact distance maps lacked stability due to the absence of semantic information. To address these limitations, we propose a novel generative model, named CoSe-CVAE, which generates contact semantic maps from object point clouds and incorporates the grasp generation pipeline for cluttered environments.
@@ II-C Grasping Evaluation in Cluttered Scenes
Concerning collision-free grasp detection in cluttered environments, extensive research has been conducted on employing neural networks to predict collision-free grasp samples from visual data, particularly in the context of two-fingered grasping setups [27, 12, 11, 43, 7, 8, 9]. However, multi-fingered hands, with their additional joints, pose greater challenges in learning implicit collision representations. Previous evaluation methods [11, 44] were often designed for a specific robotic hand and lacked the ability to generalize across different hand types. To address this complexity and identify optimal grasp candidates in cluttered environments, we develop an unified grasp evaluation model that estimate grasp scores with collision awareness.
@@ III-A Problem Statement
We utilize contact semantic map $\Omega\in\mathbb{R}^{2048\times(n+1)}$ with $2048$ points to represent the contact points between $n$ fingers of robotic hand and grasped object and points without contacts, as shown in Fig. 2 (a). Generative model CoSe-CVAE $f$ is able to estimate $N$ contact semantic maps from object point cloud $O$, as introduced in Sec. III-C. Grasp detection $F$ is presented in Sec. III-D to estimate grasp candidates based on contact semantic maps. For estimating the optimal grasp candidate $g_{\rm optimal}$ from a cluttered scene’s point cloud, grasp evaluation network $\Psi$ infers grasp qualities $q$ from partial scene point cloud $P$ and sampled hand point cloud $H$, as shown in Fig. 2 (b) and described in Sec. III-E. The partial scene point cloud is obtained through filtering the original scene points using a cylindrical region in the robotic hand’s frame, defined by a radius $r$ and height $h$, as depicted in Fig. 2 (c). The optimal grasp pose $g_{\rm optimal}$ is selected based on these inferences. The pipeline is summarized in Eq. 1.
@@ III-B Multi-Modal Multi-fingered Hand Grasping Dataset Generation
1:Input: Object database $A$, Number of sampling grasp poses $M$, Number of objects in scene $m$
2:Output: Set of multi-fingered robotic hand grasping candidates with grasp pose $g$, grasp quality $Q_{1}$, collision score $Q_{2}$, contact semantic map $\Omega$, contact distance map $\Omega_{d}$
3:// Generate dataset for single-object scenes.
4:for each object in $A$  do
5:     Estimate $g$, $Q_{1}$ based on [24], $\Omega_{d}$ based on [2].
6:     Estimate proposed $\Omega$ as detailed in III-B1.
7:end for
8:// Generate dataset for cluttered scenes.
9:for each cluttered scene do
10:     Sample $m$ objects from $A$.
11:     Construct a cluttered scene by iteratively adding objects in sampled poses. Ensure that each newly placed object does not collide with existing objects using collision detection [7].
12:     for each object in the generated cluttered scene do
13:         for each each grasp candidate of the object do
14:              Compute the collision score $Q_{2}$ between mesh of robotic hand at the candidate pose and surrounding objects using collision detection.
15:              Obtain $(g,Q_{1},Q_{2},\Omega,\Omega_{d})$.
16:         end for
17:     end for
18:end for
@@ III-B1 Contact Semantic Map Estimation
The contact semantic map is computed by estimating the nearest points on object surface to hand’s fingers. Coarse-estimated nearest points on object’s surface $P_{n}$ is calculated based on aligned distance $\epsilon(\phi_{\rm finger},O)$ [2], normal vector $n_{o}$ of object surface point and robotic hand surface point $v_{h}$, as formalized in Eq. 2. Each finger’s surface point is denoted by $\phi_{\rm finger}$ and the object point cloud is represented using $O$. The object point clouds with contact semantic labels are denoted by $P_{\rm contact}$. This process is formalized as follows:
@@ III-C Contact Semantic CVAE
Figure 4: Contact Semantic Map Generation and Grasping Detection.
@@ III-D Grasping Detection from Contact Semantic Maps
Using the generated contact semantic maps and the surface point clouds of each finger of the robotic hand, we utilize the correspondence point matching algorithm [46] to estimate the initial wrist pose. Inspired by GenDexGrasp [2] and UniGrasp [1], we optimize the wrist poses and finger tip positions by minimizing energy loss considering proposed contact semantic maps. The joint angles of the manipulator are calculated by differential inverse kinematics library mink [47]. The energy loss function $e$ is formulated as follows:
@@ III-E Grasp Evaluation Model
To identify optimal grasp in cluttered environments, we introduce an unified grasp evaluation model PointNetGPD++ to quantify grasp quality, denoted as $q$. As illustrated in Fig. 5, our network architecture integrates inputs composed of hand surface points $P$ and a partial scene point cloud $H$ surrounding the target grasped object. The partial point cloud is captured relative to the hand’s local frame. The point clouds are processed through two PointNet++ encoders, which extract latent spatial features of the hand and the scene point clouds. These latent features are concatenated and passed through a PointNet++ decoder to predict grasp classification scores. The classification includes three categories: Class 0 (grasp candidates under collision condition), Class 1 (negative grasp candidates), and Class 2 (positive grasp candidates). We employ a multi-class classification loss to guide the network’s training. Specifically, we utilize the categorical cross-entropy loss, defined as:
@@ IV-B Multi-fingered Robotic Hand Grasping Dataset
Our dataset includes 1,521 household models from our previous work [51], as well as objects from the AffordPose dataset [52]. We generated over 2,000 scenes containing cluttered scene point clouds, collision scores, grasp quality, contact semantic maps, and grasp type information, contact distance map. We predict grasp types by voting based on the contact points and the object affordance label from [52], where the grasp type is determined by the affordance with the highest number of votes. The affordance labels include handle grasping, enveloping grasp, pouring, pressing, cutting, and twisting. In this work, we did not explore grasp type information in grasp generation. Dataset examples and grasp types are shown in the Fig. 3 and Fig. 7.
@@ IV-C Model Training
Figure 8: Inference results for grasping with known and unknown hands using SOTA and our methods. (a) Estimated Contact distance map and grasp pose from GenDexGrasp [2]. (b) Contact maps and grasp pose from ContactGen [3]. (c) Contact points and grasp candidate based on UniGrasp [1]. (d)-(e) Contact semantic map and grasp candidate from our methods.
IV-D Contact Information-Guided Grasping Generation